# On anomalous dissipation induced by transport noise

### v2

@@ 1.2. Heuristics for $\nu$-dependent transport noise
Here, to motivate transport noise, we follow the heuristic motivations given in [37, Section 1.2] based on two scale arguments. Rigorous justifications of the transport noise in fluid dynamics models are given in e.g., [28, 42, 50, 65, 66].
As in [37, Section 1.2], let us assume that the vorticity field $\zeta^{\nu}$ of a turbulent fluid decomposes into large and small scales, i.e., $\zeta^{\nu}=\zeta^{\nu}_{\rm L}+\zeta_{{\rm S}}^{\nu}$ where $\zeta_{{\rm S}}^{\nu}$ and $\zeta_{\rm L}^{\nu}$ are the ‘small’ and ‘large’ scales, respectively; and
(1.8) $\partial_{t}\zeta^{\nu}_{{\rm L}}+([u_{{\rm S}}^{\nu}+u_{\rm L}^{\nu}]\cdot\nabla)\zeta_{\rm L}^{\nu}&=\nu\Delta\zeta^{\nu}_{\rm L}&\text{on }&\mathbb{T}^{2},\\ \partial_{t}\zeta_{{\rm S}}^{\nu}+([u_{{\rm S}}^{\nu}+u_{\rm L}^{\nu}]\cdot\nabla)\zeta_{{\rm S}}^{\nu}&=\nu\Delta\zeta_{{\rm S}}^{\nu}&\text{on }&\mathbb{T}^{2}.\end{aligned}\right.$
Here, $u_{\rm L}^{\nu}={\mathcal{K}}\,\zeta_{\rm L}^{\nu}$ and $u_{{\rm S}}^{\nu}={\mathcal{K}}\,\zeta_{{\rm S}}^{\nu}$ are the corresponding small and large scales velocity, respectively. Note that $\zeta^{\nu}$ solves the 2D (deterministic) NSEs. For completeness, below we also include comments on the 3D NSEs, where in (1.8) the transport term $(u^{\nu}_{*}\cdot\nabla)\zeta_{*}^{\nu}$ is replaced by $(u^{\nu}_{*}\cdot\nabla)\zeta_{*}^{\nu}-(\zeta^{\nu}_{*}\cdot\nabla)u_{*}^{\nu}$ with $*\in\{{\rm L},{\rm S}\}$. Heuristically, one could think that, in a turbulent regime, $u^{\nu}_{{\rm S}}$ varies in time very rapidly compared to $u^{\nu}_{\rm L}$. Therefore, $u^{\nu}_{{\rm S}}$ can be approximated, in time, by a white noise:
@@
where $(a_{k,\alpha})_{\alpha\in\{0,\dots,d-1\}}\subseteq{\mathbb{R}}^{d}$ is an orthonormal basis of $k^{\perp}=\{k^{\prime}\in{\mathbb{R}}^{d}\,:\,k^{\prime}\cdot k=0\}$ (this fact ensures $\nabla\cdot\sigma_{k,\alpha}=0$). Note that using (1.9) in the first of (1.8), one obtains (1.1). The choice of the Stratonovich formulation in (1.1) is due to Wong-Zakai type results which, roughly, ensures that (1.1) can be obtained as a limit of regular approximations of $\dot{W}^{k,\alpha}_{t}$. At this point, there are no physical motivations for the independence on $\nu>0$ of the coefficients $\widetilde{\theta}^{\nu}_{k}$ in (1.9). In contrast, one expects that small scales are more active at the Kolmogorov length scale $\sim\nu^{-\ell_{d}}$ where $\ell_{2}=\frac{1}{2}$ and $\ell_{3}=\frac{3}{4}$, c.f., [79, pp. 350]. As $u_{{\rm S}}^{\nu}$ in (1.9) is a model for small scales, then one expects that $k\mapsto\theta_{k}^{\nu}$ ‘accumulates’ at (high) frequencies $\sim\nu^{-\ell_{d}}$.
Due to the previous facts and the support condition in our results (1.6), we conjecture that the anomalous dissipation in Theorem 1.1, which is expected in the KB theory of 2D turbulence, holds with $\mathrm{supp}\,\theta^{\nu}\subseteq\{k\in{\mathbb{Z}}_{0}^{2}\,:\,|k|\eqsim\nu^{-3/4}\}$. In the case of passive scalars, in agreement with the fact that anomalous dissipation can be obtained by a (deterministic) Hölder continuous flow [12, 24], we instead expect that $\mathrm{supp}\,\theta^{\gamma}\subseteq\{k\in{\mathbb{Z}}_{0}^{d}\,:\,|k|\eqsim\gamma^{-\delta}\}$ for $\delta>0$ is enough for obtaining anomalous dissipation for solutions to (1.3). The proof of the above conjectures goes beyond the scope of this manuscript.
@@
$\mathcal{E}_{{\rm S}}^{\nu,\phi}(t,x)\stackrel{{\scriptstyle{\rm def}}}{{=}}\sum_{k\in{\mathbb{Z}}^{d}_{0}}\widetilde{\theta}_{k}^{\nu}\,e^{2\pi\mathrm{i}k\cdot x}\sum_{1\leq\alpha\leq d-1}a_{k,\alpha}\int_{0}^{t}\phi(s)\,\mathrm{d}W^{k,\alpha}_{s}$
@@
(1.10) $\mathbf{E}\|\mathcal{E}_{{\rm S}}^{\nu,\phi}(t,\cdot)\|_{L^{2}(\mathbb{T}^{d})}^{2}\eqsim_{d}\|\widetilde{\theta}^{\nu}\|_{\ell^{2}}^{2}\|\phi\|_{L^{2}(0,t)}^{2}\ \text{ for }t>0.$
The requirement of bounded energy of the small scales, as $\nu\downarrow 0$, therefore implies
@@
In contrast to NSEs, physical motivations for the $\gamma$-dependence of the transport noise in advection-diffusion equations (1.3) are less clear. In any case, we will use the proof of Theorem 1.2 as a guideline for the one of Theorem 1.1.
To conclude, let us stress that some criticisms can be made about the decomposition (1.8) and the corresponding ansatz (1.9). Indeed, as commented in [37, Section 1.2], the above scale separation has never been established so strictly in real fluids. However, simplified models such as (1.8)-(1.9) can be used to understand basic features of certain phenomena. The reader is referred to [63] for discussions.
@@ 1.3.1. Strategy
Here, we describe the strategy used to prove Theorem 1.2. The one of Theorem 1.1 is slightly more involved due to the presence of the convective nonlinearity. The main step in the proof of Theorem 1.2 can be roughly summarized as follows (cf., Corollary 3.5):
@@
(1.12) $\partial_{t}\varrho_{\mathrm{det}}^{\gamma}&=(\gamma+\mu)\Delta\varrho_{\mathrm{det}}^{\gamma}&\text{ on }&\mathbb{T}^{d},\\ \varrho_{\mathrm{det}}^{\gamma}(0,\cdot)&=\varrho_{0}&\text{ on }&\mathbb{T}^{d}.\end{aligned}\right.$
Here, $\approx$ means that the difference between the two quantities in (1.11) can be made as small as needed by choosing $\theta^{\gamma}$ appropriately.
@@
$2\,\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t$ =(1.13)${{=}}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}-\mathbf{E}\|\varrho^{\gamma}(1)\|_{L^{2}(\mathbb{T}^{d})}^{2}$
(1.11)${{\approx}}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}-\|\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}(\mathbb{T}^{d})}^{2}$
$\ \ \stackrel{{\scriptstyle(i)}}{{\geq}}(1-e^{-\mu})\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}$
where in $(i)$ we used that $\|\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq e^{-(\mu+\gamma)t}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq e^{-\mu t}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}$ due to the increased viscosity/diffusivity in (1.12). Now, Theorem 1.2 follows by tuning the parameter $\mu>0$ and taking the infimum over $\gamma\in(0,1)$.
@@ 1.3.2. Novelty – Improved scaling limits via Meyers’ estimates
The key ingredient in the proof of (1.11) is an improvement of a well-established scaling limit argument due to [32, 37]. Our improvement of such scaling limit argument reads roughly as follows (cf., Proposition 3.1):
@@
A sequence in $\mathcal{S}_{\ell^{2}}^{0}$ satisfying (1.14) is given in [61, eq. (1.9)] (see also (3.4)).
The main novelty and improvement compared to the above-mentioned works of (1.15) is that the limit is established in a space of zero smoothness uniformly in time. More precisely, in previous works, limits as in (1.15) were only established with $L^{2}(\mathbb{T}^{d})$ replaced by $H^{-\varepsilon}(\mathbb{T}^{d})$ for some $\varepsilon>0$ (cf., [32, Proposition 3.7] and [37, Theorem 1.4]). Let us stress that such improvement is of central importance to establish (1.11) and thus for our approach to anomalous dissipation. Indeed, (1.15) and energy estimates readily imply (1.11) for a sufficiently large $n\geq 1$; while this is not true if one uses $H^{-\varepsilon}(\mathbb{T}^{d})$ instead of $L^{2}(\mathbb{T}^{d})$ in (1.15). Interestingly, we can also prove that (1.15) holds with $L^{2}(\mathbb{T}^{d})$ replaced by $H^{r}(\mathbb{T}^{d})$ for some $r(\gamma,\mu)>0$. The latter fact is needed to deal with 2D NSEs, see the proof of Corollary 4.4.
The key tool behind (1.15) are the stochastic Meyers’ estimates proven in Section A. In particular, they imply the existence of $r_{0}(\gamma,\mu)>0$ for which
@@
(1.17) $\sup_{n\geq 1}\|(\theta^{n}_{k}\sigma_{k,\alpha})_{k,\alpha}\|_{L^{\infty}(\mathbb{T}^{d};\ell^{2})}<\infty\ \ \text{ while }\ \ \sup_{n\geq 1}\|(\theta^{n}_{k}\sigma_{k,\alpha})_{k,\alpha}\|_{H^{r}(\mathbb{T}^{d};\ell^{2})}=\infty.$
Let us stress that the supremum over $n\geq 1$ in the above is essential it can hold that #$\#\{k\,:\,\theta_{k}^{n}\neq 0\}<\infty$ and therefore $(\theta^{n}_{k}\sigma_{k,\alpha})_{k,\alpha}\in C^{\infty}(\mathbb{T}^{d};\ell^{2})$ for each fixed $n\geq 1$. In particular, (1.16) cannot be derived from well-known results on $L^{p}$-theory of SPDEs, as it always requires some degree of regularity of the coefficients and provides a high improvement of the regularity of solutions the SPDE under consideration, see e.g., [8, 57]. In contrast, Meyers’ estimates provide a small improvement on the regularity of solutions of (S)PDEs with bounded, measurable and parabolic coefficients; which is exactly the setting one faces up with when dealing with the scaling limit (1.14)-(1.15). Indeed, the boundedness is given by (1.17), while the parabolicity is uniform due to the Stratonovich formulation of the transport noise, see the proof of Lemma 3.2 for details. To conclude, let us stress that, due to the lack of uniform smoothness as described in (1.17), (1.16) cannot hold with $r_{0}$ large. Counterexamples in the deterministic counterpart can be found in [14, 67].
@@ 1.3.3. Physical interpretation – Connection with homogenization
Arguing as in [1, Subsection 2.2], the scaling limit result of (1.14)-(1.15) can be interpreted as an homogenization result for the SPDE (1.3). Indeed, as discussed in Subsection 1.2, the transport noise term in (1.3) (or, more precisely, for 2D NSEs (1.1) for which a similar discussion applies) can be interpreted as the contribution of the ‘small scales’ of the fluid flow. Therefore, reasoning as in [1, Subsection 2.2], the limit (1.14) can be interpreted as ‘zooming out’ from small scales and the PDE (1.12) can be thought of as the ‘effective equation’ for the SPDE (1.3). In our context, the ‘microscopic’ parameter is $\varepsilon=\|\theta^{n}\|_{\ell^{\infty}}\downarrow 0$, cf., (1.14).
The homogenization viewpoint is also interesting from a mathematical perspective. Indeed, the main tool to achieve (1.15) are the (stochastic) Meyers’ estimates which also play a fundamental role in the context of homogenization of PDEs (see e.g., [10, 11, 46, 75]) since, as it is well-known in homogenization, the only quantities remaining uniform along the process of zooming out from small scales (i.e., $\varepsilon\downarrow 0$) are measurability, boundedness and parabolicity.

### v3

@@ 1.2. Heuristics for $\nu$-dependent transport noise
Here, to motivate transport noise, we follow the heuristic motivations given in [42, Section 1.2] based on two scale arguments. Rigorous justifications of the transport noise in fluid dynamics models are given in, e.g., [31, 47, 56, 72, 73].
As in [42, Subsection 1.2], let us assume that the vorticity field $\zeta^{\nu}$ of a turbulent fluid decomposes into large and small scales, i.e., $\zeta^{\nu}=\zeta^{\nu}_{\rm L}+\zeta_{{\rm S}}^{\nu}$ where $\zeta_{{\rm S}}^{\nu}$ and $\zeta_{\rm L}^{\nu}$ are the ‘small’ and ‘large’ scales, respectively; and
(1.10) $\left\{\begin{aligned} \partial_{t}\zeta^{\nu}_{{\rm L}}+([u_{{\rm S}}^{\nu}+u_{\rm L}^{\nu}]\cdot\nabla)\zeta_{\rm L}^{\nu}&=\nu\Delta\zeta^{\nu}_{\rm L}&\text{on }&\mathbb{T}^{2},\\ \partial_{t}\zeta_{{\rm S}}^{\nu}+([u_{{\rm S}}^{\nu}+u_{\rm L}^{\nu}]\cdot\nabla)\zeta_{{\rm S}}^{\nu}&=\nu\Delta\zeta_{{\rm S}}^{\nu}&\text{on }&\mathbb{T}^{2}.\end{aligned}\right$.
Here, $u_{\rm L}^{\nu}={\mathcal{K}}\,\zeta_{\rm L}^{\nu}$ and $u_{{\rm S}}^{\nu}={\mathcal{K}}\,\zeta_{{\rm S}}^{\nu}$ are the corresponding small and large scales velocity, respectively. Note that $\zeta^{\nu}$ solves the 2D (deterministic) NSEs. For completeness, below we also include comments on the 3D NSEs, where in (1.10) the transport term $([u^{\nu}_{{\rm S}}+u^{\nu}_{{\rm L}}]\cdot\nabla)\zeta_{*}^{\nu}$ is replaced by $([u^{\nu}_{{\rm S}}+u^{\nu}_{{\rm L}}]\cdot\nabla)\zeta_{*}^{\nu}-(\zeta^{\nu}_{*}\cdot\nabla)[u^{\nu}_{{\rm S}}+u^{\nu}_{{\rm L}}]$ with $\in\{{\rm L},{\rm S}\}$, cf. again [42, Subsection 1.2]. Heuristically, one could think that, in a turbulent regime, $u^{\nu}_{{\rm S}}$ varies in time very rapidly compared to $u^{\nu}_{\rm L}$. Therefore, $u^{\nu}_{{\rm S}}$ can be approximated, in time, by a white noise:
@@
where $(a_{k,\alpha})_{\alpha\in\{0,\dots,d-1\}}\subseteq{\mathbb{R}}^{d}$ is an orthonormal basis of $k^{\perp}=\{k^{\prime}\in{\mathbb{R}}^{d}\,:\,k^{\prime}\cdot k=0\}$ (this fact ensures $\nabla\cdot\sigma_{k,\alpha}=0$). Note that using (1.11) in the first of (1.10), one obtains (1.1). The choice of the Stratonovich formulation in (1.1) is due to Wong-Zakai-type results, which, roughly, ensure that (1.1) can be obtained as a limit of regular approximations of $\dot{W}^{k,\alpha}_{t}$. At this point, there are no physical motivations for the independence of the coefficients $\widetilde{\theta}^{\nu}_{k}$ on $\nu>0$ in (1.11). In contrast, one expects that small scales are more active at the Kolmogorov length scale $\sim\nu^{-\ell_{d}}$ where $\ell_{2}=\frac{1}{2}$ and $\ell_{3}=\frac{3}{4}$, c.f., [87, pp. 350]. As $u_{{\rm S}}^{\nu}$ in (1.11) is a model for small scales, then one expects that $k\mapsto\theta_{k}^{\nu}$ ‘accumulates’ at (high) frequencies $\sim\nu^{-\ell_{d}}$.
Due to the previous facts and the support condition in our results (1.7), we conjecture that the anomalous dissipation in Theorem 1.1, which is expected in the KB theory of 2D turbulence, holds if and only if $\mathrm{supp}\,\theta^{\nu}\subseteq\{k\in{\mathbb{Z}}_{0}^{2}\,:\,|k|\eqsim\nu^{-\delta}\}$ for some $\delta\geq\frac{1}{2}$ independent of $\nu$. In the case of passive scalars, in agreement with the fact that anomalous dissipation can be obtained by a (deterministic) Hölder continuous flow [14, 27], we instead expect that $\mathrm{supp}\,\theta^{\gamma}\subseteq\{k\in{\mathbb{Z}}_{0}^{d}\,:\,|k|\eqsim\gamma^{-\delta}\}$ for $\delta>0$ is sufficient for obtaining anomalous dissipation for solutions to (1.3). The proof of the above conjectures goes beyond the scope of this manuscript.
@@
$\mathcal{E}_{{\rm S}}^{\nu,\phi}(t,x)\stackrel{{\scriptstyle{\rm def}}}{{=}}\Big|\sum_{k\in{\mathbb{Z}}^{d}_{0}}\widetilde{\theta}_{k}^{\nu}\,e^{2\pi\mathrm{i}k\cdot x}\sum_{1\leq\alpha\leq d-1}a_{k,\alpha}\int_{0}^{t}\phi(s)\,\mathrm{d}W^{k,\alpha}_{s}\Big|^{2}$
@@
(1.12) $\mathbf{E}\int_{\mathbb{T}^{d}}\mathcal{E}_{{\rm S}}^{\nu,\phi}(t,x)\,\mathrm{d}x\eqsim_{d}\|\widetilde{\theta}^{\nu}\|_{\ell^{2}}^{2}\|\phi\|_{L^{2}(0,t)}^{2}\ \text{ for }t>0.$
The requirement that the energy of the small scales is bounded, as $\nu\downarrow 0$, therefore implies
@@
To conclude, let us stress that some criticisms can be made about the decomposition (1.10) and the corresponding ansatz (1.11). Indeed, as commented in [42, Section 1.2], the above scale separation has never been established so strictly in real fluids. However, simplified models such as (1.10)-(1.11) can be used to understand basic features of certain phenomena. The reader is referred to [69] for discussions.
@@ 1.3.1. Strategy
Here, we describe the strategy used to prove Theorem 1.2. The proof of Theorem 1.1 is slightly more involved due to the presence of the convective nonlinearity. The main step in the proof of Theorem 1.2 can be roughly summarized as follows (cf. Corollary 3.5):
@@
(1.14) $\left\{\begin{aligned} \partial_{t}\varrho_{\mathrm{det}}^{\gamma}&=(\gamma+\mu)\Delta\varrho_{\mathrm{det}}^{\gamma}&\text{ on }&\mathbb{T}^{d},\\ \varrho_{\mathrm{det}}^{\gamma}(0,\cdot)&=\varrho_{0}&\text{ on }&\mathbb{T}^{d}.\end{aligned}\right$.
Here, $\approx$ means that the difference between the two quantities in (1.13) can be made as small as needed by choosing $\theta^{\gamma}$ appropriately. Below, we use that $\varrho_{0}$ has mean zero by assumption, see Theorem 1.2.
@@
[…]
$\ \ \stackrel{{\scriptstyle(i)}}{{\geq}}(1-e^{-\mu/(4\pi^{2})})\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}$
where in $(i)$ we used that from the energy inequality for (1.14) it holds that $\|\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq e^{-[(\mu+\gamma)t]/(4\pi^{2})}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq e^{-(\mu t)/(4\pi^{2})}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}$, due to the increased viscosity/diffusivity in (1.14). Here, we used $\int_{\mathbb{T}^{d}}\varrho_{\mathrm{det}}^{\gamma}(t,\cdot)\,\mathrm{d}x=0$ for all $t\geq 0$ as $\int_{\mathbb{T}^{d}}\varrho_{0}\,\mathrm{d}x=0$, and that the Poincaré constant is $1/(2\pi)$ (the latter follows from the Plancherel inequality and $\mathbb{T}={\mathbb{R}}/{\mathbb{Z}}$). Hence, Theorem 1.2 follows by taking the $\liminf_{\gamma\downarrow 0}$ in the above.
@@ 1.3.2. Novelty – Improved scaling limits via Meyers’ estimates
The key ingredient in the proof of (1.13) is an improvement of a well-established scaling limit argument due to [36, 42]. Our improvement of such a scaling limit argument reads roughly as follows (cf. Proposition 3.1):
@@
A sequence in $\mathcal{S}_{\ell^{2}}^{0}$ satisfying (1.16) is given in [67, eq. (1.9)] (see also (3.4)).
The main novelty and improvement compared to the above-mentioned works of (1.17) is that the limit is established in a space of zero smoothness uniformly in time. More precisely, in previous works, limits as in (1.17) were only established with $L^{2}(\mathbb{T}^{d})$ replaced by $H^{-\varepsilon}(\mathbb{T}^{d})$ for some $\varepsilon>0$ (cf. [36, Proposition 3.7] and [42, Theorem 1.4]). The validity of (1.17) requires (1.3) to be parabolic, i.e., $\gamma>0$, and discriminates between the parabolic ($\gamma>0$) and hyperbolic ($\gamma=0$) regimes; see [25, 37] for some results in the hyperbolic setting. It is worth noticing that (1.17) implies the non-convergence of $\nabla\varrho^{n}$ to $\nabla\varrho_{\mathrm{det}}^{\gamma}$, see Remark 3.4.
The improvement in (1.17) is of central importance to establish (1.13), and thus for our approach to anomalous dissipation. Indeed, (1.17) and energy estimates readily imply (1.13) for a sufficiently large $n\geq 1$; while this is not true if one uses $H^{-\varepsilon}(\mathbb{T}^{d})$ instead of $L^{2}(\mathbb{T}^{d})$ in (1.17). Interestingly, we can also prove that (1.17) holds with $L^{2}(\mathbb{T}^{d})$ replaced by $H^{r}(\mathbb{T}^{d})$ for some $r(\gamma,\mu)>0$. The latter fact is needed to deal with, e.g., 2D NSEs (1.1) (see the proof of Corollary 4.4) and in the velocity formulation in Section 5.
The key tool behind (1.17) is the stochastic Meyers’ estimates proven in Section A. In particular, they imply the existence of $r_{0}(\gamma,\mu)>0$ for which
@@
(1.19) $\sup_{n\geq 1}\|(\theta^{n}_{k}\sigma_{k,\alpha})_{k,\alpha}\|_{L^{\infty}(\mathbb{T}^{d};\ell^{2})}<\infty\ \ \text{ while }\ \ \sup_{n\geq 1}\|(\theta^{n}_{k}\sigma_{k,\alpha})_{k,\alpha}\|_{H^{r}(\mathbb{T}^{d};\ell^{2})}=\infty;$
see [3, Proposition 4.1] for details. Let us stress that the supremum over $n\geq 1$ in the above is essential as it can hold that $\{k\,:\,\theta_{k}^{n}\neq 0\}<\infty$ and therefore $(\theta^{n}_{k}\sigma_{k,\alpha})_{k,\alpha}\in C^{\infty}(\mathbb{T}^{d};\ell^{2})$ for each fixed $n\geq 1$. In particular, (1.18) cannot be derived from well-known results on $L^{p}$-theory of SPDEs, as it always requires some degree of regularity of the coefficients and provides a high improvement of the regularity of solutions to the SPDE under consideration, see e.g., [9, 63]. In contrast, Meyers’ estimates provide a small improvement on the regularity of solutions of (S)PDEs with bounded, measurable, and parabolic coefficients, which is exactly the setting one encounters when dealing with the scaling limit (1.16)-(1.17). Indeed, the boundedness is given by (1.19), while the parabolicity is uniform due to the Stratonovich formulation of the transport noise, see the proof of Lemma 3.2 for details. To conclude, let us stress that, due to the lack of uniform smoothness as described in (1.19), (1.18) cannot hold with $r_{0}$ large. Counterexamples in the deterministic counterpart can be found in [16, 74].
@@ 1.3.3. Physical interpretation – Connection with homogenization
Arguing as in [1, Subsection 2.2], the scaling limit result of (1.16)-(1.17) can be interpreted as a homogenization result for the SPDE (1.3). Indeed, as discussed in Subsection 1.2, the transport noise term in (1.3) (or, more precisely, for 2D NSEs (1.1) for which a similar discussion applies) can be interpreted as the contribution of the ‘small scales’ of the fluid flow. Therefore, reasoning as in [1, Subsection 2.2], the limit (1.16) can be interpreted as ‘zooming out’ from small scales, and the PDE (1.14) can be thought of as the ‘effective equation’ for the SPDE (1.3). In our context, the ‘microscopic’ parameter is $\varepsilon=\|\theta^{n}\|_{\ell^{\infty}}\downarrow 0$, cf. (1.16).
The homogenization viewpoint is also interesting from a mathematical perspective. Indeed, the main tool to achieve (1.17) are the (stochastic) Meyers’ estimates which also play a fundamental role in the context of homogenization of PDEs (see e.g., [12, 13, 51, 82]) since, as it is well-known in homogenization, the only quantities remaining uniform along the process of zooming out from small scales (i.e., $\varepsilon\downarrow 0$) are measurability, boundedness, and parabolicity.